OpenStudy (anonymous):

FAN AND MEDAL!!! If (x + 3)2 = 42, what is the approximate positive value of x? 0.48 3.48 5.48 6.48

OpenStudy (yanasidlinskiy):

Plug in the answer choices and see which one comes out the closest.

OpenStudy (anonymous):

so its asking what is the value of X+?. try each answer and see which one works!

OpenStudy (anonymous):

i have and i keep getting some weird answer

OpenStudy (anonymous):

is d question correct????

OpenStudy (bibby):

better yet, isolate x by squaring both sides, than solve for x you need to use a calculator though

OpenStudy (anonymous):

okay well 0.48+3*2=? 3.48+3*2=? 5.48+3*2=? 6.48+3*2+?

OpenStudy (anonymous):

... so d? i tried that and got a weird answer

OpenStudy (anonymous):

oh your right they all come out different. hmmm let me look at it a minute and ill see what i can find out.

OpenStudy (anonymous):

thanks @jenks1214

OpenStudy (anonymous):

hey isn't x=18.......

OpenStudy (anonymous):

I think the answer is B.

OpenStudy (anonymous):

can you tell me how you got that ?? @ashhhrodriguez

OpenStudy (anonymous):

yeah thats what i though too @srev98

OpenStudy (bibby):

\((x + 3)^2 = 42\\(x+3)=\sqrt{42}\\x+3=6.48074069841\)

OpenStudy (anonymous):

Instead of it being times two. I think its exponent of 2. So 3.48+3=6.48. 6.48x6.48= 41.99

OpenStudy (anonymous):

wait so is it b or d ?? @bibby @ashhhrodriguez

OpenStudy (anonymous):

you guys just confused me

OpenStudy (bibby):

you're not even reading what we're saying

OpenStudy (anonymous):

nvm i got it its b

OpenStudy (anonymous):

@bibby wrote x+3=6.48074 You have to subtract 3 from both sides getting 3.48

OpenStudy (anonymous):

yea I know my sister just showed me how to do it

OpenStudy (anonymous):

lol ok yeah

OpenStudy (anonymous):

haha yea